How many grams of (NH4)2SO4 (fw = 132.14 g/mol) should be added to 425 mL of a 0.258 M solution of NH3 to achieve a buffer solution with a pH of 9.44? Kb for NH3 = 1.8 x 10-5

1
Expert's answer
2021-11-17T05:01:01-0500

Molarity of NH3 solution = moles of NH3/ volume of solution

moles of "NH_3=0.2\u00d70.5=0.1"

"pOH=14\u2212pH"

"pOH=14\u22129.44=4.56"

"pK_b\n\n\u200b=\u2212log[K_b\n\n\u200b]=\u2212log[1.78\u00d710^{\u22125}\n\n]=4.74"

"pOH=pK_b\n\n\u200b\u2212log\\frac{\n\n\n\n[base]}{[salt]}"

"log\\frac{\n\n\n\n[base]}{[salt]}\n\n\u200b=4.74\u22124.56=0.18"

"\\frac{[base]}{[salt]}\n\n\u200b=antilog[0.18]=1.51"


"[base] =" moles of "NH_3\n\n\u200b=0.1"


"[salt]=\\frac{\n\n\n\n0.1}{1.51}\n\n\u200b=0.07"

Moles of "(NH_4\n\n\u200b)^+\n\n=0.07"

"(NH_4\n\n\u200b)_2\n\n\u200bSO_4\n\n\u200b\u21cc2NH_4\n\n\u200b^+\n\n+(SO_4\n\n\u200b)^{\u22122}"


moles of "(NH_4\n\n\u200b)_2\n\n\u200bSO_4\n\n\u200b=\\frac{0.07}{2}=0.035"


amount of "(NH4\n\n\u200b)2\n\n\u200bSO_4\n\n\u200b=0.035\u00d7132.14=4.6249g"